# Estimating Unknown Population Sizes Using the Hypergeometric Distribution

### v1

@@ Abstract
The multivariate hypergeometric distribution describes sampling without replacement from a discrete population of elements divided into multiple categories. Addressing a gap in the literature, we tackle the challenge of estimating discrete distributions when both the total population size and the sizes of its constituent categories are unknown. Here, we propose a novel solution using the hypergeometric likelihood to solve this estimation challenge, even in the presence of severe under-sampling. We develop our approach to account for a data generating process where the ground-truth is a mixture of distributions conditional on a continuous latent variable, such as with collaborative filtering, using the variational autoencoder framework. Empirical data simulation demonstrates that our method outperforms other likelihood functions used to model count data, both in terms of accuracy of population size estimate and in its ability to learn an informative latent space. We demonstrate our method’s versatility through applications in NLP, by inferring and estimating the complexity of latent vocabularies in text excerpts, and in biology, by accurately recovering the true number of gene transcripts from sparse single-cell genomics data.
@@ 1 Introduction
The classic Pólya urn model (Eggenberger & Pólya, 1923) describes the process of randomly sampling from an urn containing balls of various colors, and is used to illustrate common discrete probability distributions that form the core building block of many probabilistic machine learning models. When balls are sampled from the urn with replacement, the distribution of counts of balls of each color is described by the multinomial distribution, whereas the hypergeometric distribution describes sampling without replacement. The hypergeometric distribution becomes important for successful modelling when the selection of one element from the distribution affects the probabilities of selecting subsequent elements (i.e. the counts are not independent), and when the number of elements sampled is significant compared to the population size. The hypergeometric distribution also enables the direct modeling of category counts, as opposed to category probabilities, which is often of interest for downstream analyses.
@@
There are many settings where it is valuable to model count data directly and to capture the dependence between category counts. For example, in the context of collaborative filtering, song or movie play counts and shopping basket item counts can be thought of as being sampled from a larger catalogue. Likewise, a text document can be viewed as a bag-of-words that is sampled from a larger, underlying bag-of-words. In all of these cases the assumption of sampling without replacement is justifiable because the magnitude of counts is relatively small, and there is dependence between counts. For example, watching a movie changes the probability of it being watched a second time, adding an item to a virtual shopping basket changes the probability of it being added again, and selecting one word changes the probability of its synonyms and related terms also appearing.
@@
This setting is even more common in biology due to the experimental sampling methods used. In the field of single-cell genomics, the quantity of gene transcript count data measured at the resolution of individual cells is accumulating at an exponential rate. As we show in Section 4.4, high-throughput experiments inherently sample without replacement, and with careful interpretation this single-cell data promises to answer important questions in biology and health. Due to the frequent occurrence of finite, discrete populations in nature, there are many other natural phenomena that can be modelled using the hypergeometric distribution.
@@
Furthermore, the aforementioned applications are typically characterized by intrinsic low-rank structure. Movie and music choices are driven by an underlying set of preferences, words in a text passage are associated with the document’s topic, and a cell’s gene transcript counts depend strongly on its cell type. This latent structure induces correlation between features, such as songs of the same genre, and suggests that a latent-factor model is appropriate.
@@
Despite the fundamental importance of the hypergeometric distribution - it being one of the few probability distributions for modeling count data - there is currently no effective way of estimating its parameters in many common settings of interest, including in the presence of high-dimensional data with intrinsic low-rank structure. In this paper, we present a simple yet powerful method to solve an as yet unaddressed core problem for the hypergeometric distribution, namely inferring a mixture of discrete distributions where both the size of the total population and its constituent categories are unknown. We first show that parameter inference is tractable when there are two or more categories using empirical simulations, and that our method outperforms existing likelihood-based approaches. To demonstrate its value in real-world settings, we use our method to infer a latent vocabulary for reading passages that can be used to assess their complexity. Finally, we address the technical limitations of genomics experimental methods and recover missing gene transcript counts in high-throughput single-cell data.
Our novel contributions are the proposed method, the first empirical demonstration that this problem is tractable, and its successful application to two real-world problems.
@@ 2.1 The Hypergeometric Distribution
Consider an urn that contains $N$ balls divided into $K=2$ categories (e.g. colors): $N_{1}$ balls are white and $N_{2}=N-N_{1}$ are black. Each ball has equal probability of being selected from the urn. If we sample $n<N$ balls without replacement, obtaining $c_{1}$ white counts and $c_{2}$ black counts, the distribution of counts for the number of sampled balls of each color is given by the hypergeometric distribution, whose joint probability mass function is (Moivre, 1711):
@@
Typically $N$ is assumed to be known, however here we consider the setting where neither $N$ nor $N_{1}$ or $N_{2}$ are known. In the typical setting the distribution for $K=2$ is called the univariate hypergeometric distribution, with $K>2$ referred to as multivariate, but because in our problem setting there are already two unknown variables when $K=2$, we use the name hypergeometric distribution for any $K\geq 2$.
@@ 2.2 Existing Maximum Likelihood Estimators
Known total population size: When the total population size $N$ is known, the object is to estimate the true number of elements $N_{i}$ in each constituent category $i\in 1,\ldots,K$. The maximum likelihood estimator is then essentially the known population size scaled by the sample frequency of each category, with adjustment to ensure a correct integer solution (Oberhofer & Kaufmann, 1987).
Unknown total population size: In the more complex case, known as the capture-recapture problem (Darroch, 1958), the total population size is unknown. To estimate the total population, a sample is first drawn from the underlying distribution. All objects belonging to one of the categories in this sample are tagged, and all sampled elements are returned to the population. A second sample is then taken, and the number of tagged samples that reappear permits the estimation of the total population size using maximum likelihood. This method has found important applications in biology and ecology. However, it depends on the ability to tag and resample the same population, which is often not possible in practice.
Additionally, Tohma et al. (1991) propose a number of options for approximating the estimation of the parameters of the hypergeometric distribution when the category sizes are unknown, including an approximation using the likelihood, however they do not maximize the likelihood directly.
@@ 2.3 Related Methods
The multinomial distribution, which is the limiting form of the hypergeometric distribution when the sample size is negligible relative to the population ($N>>n$), appears frequently in machine learning literature. For example, it is often used to model counts indirectly by instead using the relative frequency of counts. Latent Dirichlet allocation (Blei et al., 2003) uses a hierarchical generative framework to model the distribution of topics, documents over topics, and word (category) counts from a vocabulary (population) over documents. The distributions over topics and words are multinomial, with word counts being transformed into their relative frequency in the document. (Liang et al., 2018) use a variational autoencoder with multinomial likelihood applied to collaborative filtering for recommender systems. The click data and play counts are binarized to accommodate the multinomial likelihood.
@@ 3.1 Single Ground-Truth Distribution
We first assume that there exists a single true discrete population of elements, with total number of elements $N$ divided into $K$ categories. We consider the setting where none of the category sizes are known, and we wish to estimate them from the data; that is, we do not know any category size $N_{i}$, nor the total population size $N=\sum_{i=1}^{K}N_{i}$. Because we aim to address real-world problems, we assume severe under-sampling of the underlying distribution ($n<<N$).
@@
$P({\bm{c}}_{1},\ldots,{\bm{c}}_{T}|N_{1},\ldots,,N_{K})=\prod_{t=1}^{T}\frac{\prod_{i=1}^{K}\binom{N_{i}}{c_{t,i}}}{\binom{\sum_{i=1}^{K}N_{i}}{\sum_{i=1}^{K}c_{t,i}}}$ (3)
@@
$\begin{split}&\log P({\bm{c}}_{1},\ldots,{\bm{c}}_{T}|N_{1},\ldots,,N_{K})=\\ &\sum_{t=1}^{T}\left[\sum_{i=1}^{K}\log\binom{N_{i}}{c_{t,i}}-\log\binom{\sum_{i=1}^{K}N_{i}}{\sum_{i=1}^{K}c_{t,i}}\right]\end{split}$ (4)
The hypergeometric distribution is not part of the exponential family, so it is not clear that a closed-form maximum likelihood estimator exists, therefore we turn to numerical optimization methods. To enable continuous optimization, we consider a continuous and differentiable relaxation of the log-likelihood by replacing the factorials in the binomial coefficient with the gamma function, which is the extension of the factorial to real-numbered arguments.
@@
Using this log-likelihood for the parameter set $\theta=\{N_{1},\ldots,N_{K}\}$, we perform maximum likelihood estimation to obtain the MLE $\hat{\theta}\equiv\{\hat{N}_{1},\ldots,\hat{N}_{K}\}=\operatorname*{arg\,max}_{\theta}\log p({\bm{c}}_{1},\ldots,{\bm{c}}_{T}|\theta)$.
The binomial coefficient is not defined if $c_{k}>\hat{N}_{k}$, which would correspond to the impossible scenario of sampling more balls of a given color than are present in the urn. To impose the requirement $N_{k}\geq c_{k}$, we add a violation penalty $C_{\text{viol}}$ to the negative log-likelihood we seek to minimize, and we threshold any estimates of $\hat{N}_{k}<c_{k}$ at $c_{k}$ before evaluating the likelihood. We threshold at the observed sample value $c_{i}$ as opposed to the minimum across all samples $\min_{i}c_{i}$ to remain as general as possible. This is because in the case of a mixture of distributions we do not know which observation originates from which underlying distribution, and hence do not know what the correct minimum is.
$C_{\text{viol}}$ $=\sum_{i=1}^{K}\min(0,c_{i}-\hat{N}_{i})$ (6)
@@
In this paper we are interested in modelling scenarios where we have access to many samples drawn from the same underlying population, but where this population is under-sampled. Specifically, if $N$ is the total population size, we assume that we observe samples with at most $n_{max}$ objects drawn from the ground-truth distribution in each trial, giving a sample fraction $f_{max}=n_{max}/N$. For example, if $f_{max}=0.5$ and the true distribution has 100 elements, no observation will contain more than 50 samples.
@@ 3.2 Mixture of Ground-Truth Distributions
While the hypergeometric likelihood can be used to directly estimate the ground truth for a single underlying distribution (Section 4.1) using maximum likelihood, datasets generated by real world processes are often better represented by a mixture of distributions. Therefore, we next propose a generative latent variable model $p({\bm{c}},{\bm{z}})=p({\bm{c}}|{\bm{z}})p({\bm{z}})$ to jointly model the observed counts conditional on a continuous latent variable.
We extend our method to allow for the estimation of a high-dimensional distribution that is conditional on a continuous latent variable, allowing it to capture a continuous mixture of count distributions. This modelling assumption is essential when the true distribution is considerably undersampled, leading to sparsity, and so information needs to be shared between similar observations to successfully model the data. The variational autoencoder (VAE) (Kingma & Welling, 2014) is a powerful framework for performing efficient estimation of the generative distribution parameters in the presence of a continuous latent variable and big data.
@@
$\begin{split}\mathcal{L}(\psi,\phi;{\bm{c}}_{t})=&-D_{\text{KL}}(q_{\phi}({\bm{z}}|{\bm{c}}_{t})||p({\bm{z}}))+\log p_{\psi}({\bm{c}}_{t}|{\bm{z}})\\ &+\sum_{i=1}^{K}\max(0,c_{i}-\hat{N}_{i})\end{split}$ (8)
The first term is the KL divergence between the true and approximate posterior, the second term is the log-likelihood, and the third term is the violation penalty. Optimization of this objective enables the estimation of counts under-sampled from a mixture of distributions.
@@ 4.1 Tractability of Inferring Ground-Truth Distributions
Using empirical data simulation, we begin by demonstrating that this inference problem is tractable in the ideal case where observations are sampled from a single discrete ground-truth distribution. Despite its simplicity, this scenario has yet to be shown tractable in the literature, and in fact appears to have generally been assumed intractable.
@@
To demonstrate the typical behavior, we show results for a simulation for the scenario where $N_{1}=70$ and $N_{2}=30$ ($N=100$, $K=2$, $f_{max}=0.4$), evaluating the negative log-likelihood (NLL) for all possible combinations of $\hat{N}_{1}$ and $\hat{N}_{2}$. Figure 1 shows the resulting loss landscape, and we see that the minimum NLL value corresponds to the true $N_{1}$ and $N_{2}$ values. The lowest NLL region is elongated along the line of correct $N_{2}/N_{1}$ ratio. We note that the difference in NLL between the minimum and surrounding estimates is small, especially for estimates with the same ratio $N_{2}/N_{1}$.
@@
We next investigate the robustness of the maximum likelihood estimate for a wide range of numbers of observations $T$ and max sample fractions $f_{max}$. As the data is discrete, we use the Manhattan distance between the true and estimated counts to quantify the estimation error. In Figure 2, we see that the location of the minimum NLL converges to the true value as the number of observations increases, for different levels of under-sampling. As expected, a higher max sample fraction results in faster convergence and more accurate estimates, as quantified by lower error for the same number of samples. We extend this experiment to $K=3$ (Figure 7, Appendix E), and find that increasing the number of categories reduces the number of observations required to reach zero error. Here we use a confidence interval of 50% for visual clarity, and the same figure with a 90% confidence interval is included in Appendix E.
Figure 2: Maximum likelihood estimate Manhattan error for different numbers of observations at different max sample fractions. 50% confidence interval over 50 random seeds. $K=2,N=100$ $(N_{1}=40,N_{2}=60)$

### v2

@@ Abstract
The multivariate hypergeometric distribution describes sampling without replacement from a discrete population of elements divided into multiple categories. Addressing a gap in the literature, we tackle the challenge of estimating discrete distributions when both the total population size and the category sizes are unknown. Here, we propose a novel solution using the hypergeometric likelihood to solve this estimation problem, even in the presence of severe under-sampling. Our approach accounts for a data generating process where the ground-truth is a mixture of distributions conditional on a continuous latent variable, as seen in collaborative filtering, using the variational autoencoder framework. Empirical data simulation demonstrates that our method outperforms other likelihood functions used to model count data, both in terms of accuracy of population size estimate and learning an informative latent space. We showcase our method’s versatility through applications in NLP, by inferring and estimating the complexity of latent vocabularies in reading passage excerpts, and in biology, by accurately recovering the true number of gene transcripts from sparse single-cell genomics data.
Machine Learning, ICML
@@ 1 Introduction
The classic Pólya urn model (Eggenberger & Pólya, 1923) describes the process of randomly sampling from an urn containing balls of various colors, and is used to illustrate common discrete probability distributions that form the core building block of many probabilistic machine learning models. When balls are sampled from the urn with replacement, the distribution of counts of balls of each color is described by the multinomial distribution, whereas the hypergeometric distribution describes sampling without replacement. The hypergeometric distribution becomes important for successful modelling when the selection of one element from the distribution affects the probabilities of selecting subsequent elements (i.e. the counts within as sample are not independent), and when the number of elements sampled is significant compared to the population size. The hypergeometric distribution also enables the direct inference of category counts, as opposed to category probabilities, which is often of interest for downstream analyses and interpretability.
@@
There are many settings where it is valuable to model count data directly and to capture the dependence between category counts. For example, in the context of collaborative filtering, song or movie play counts (Van den Oord et al., 2013) and shopping basket item counts (Faggioli et al., 2020) can be thought of as being sampled from a larger catalogue. Likewise, a text document can be viewed as a bag-of-words that is sampled from a larger, underlying bag-of-words (Goldberg, 2022). In all of these cases the assumption of sampling without replacement is justifiable because the magnitude of counts is relatively small, and there is dependence between these counts. For example, watching a movie changes the probability of it being watched a second time, adding an item to a virtual shopping basket changes the probability of it being added again, and selecting one word changes the probability of its synonyms and related terms also appearing in a text.
@@
This setting is even more common in biology due to the experimental sampling methods used (DePatta Pillar, 1998). In the field of single-cell genomics, the quantity of gene transcript count data measured at the resolution of individual cells is accumulating at an exponential rate (Svensson et al., 2018). As we show in Section 4.4, high-throughput experiments inherently sample without replacement, and with careful interpretation this single-cell data promises to answer important questions in biology and health. In the field of microbiome research, another potential application is the estimation of microbial population sizes and subsequent differential abundance analysis (Morton et al., 2019a, b). Due to the frequent occurrence of finite, discrete populations in nature, there are many other natural phenomena that can be modelled using the hypergeometric distribution (Holmes & Huber, 2018).
@@
Furthermore, the aforementioned applications are typically characterized by intrinsic low-rank structure. Movie and music choices are driven by an underlying set of preferences (Feuerverger et al., 2012), words in a text passage are associated with the document’s topic (Vayansky & Kumar, 2020), and a cell’s gene transcript counts depend strongly on its cell type (Grønbech et al., 2020). This latent structure induces correlation between features, such as songs of the same genre, and suggests that a latent-factor model is appropriate.
@@
Despite the fundamental importance of the hypergeometric distribution - it being one of the few fundamental probability distributions for modeling sampling - there is currently no effective way of inferring its parameters in many common settings of interest, including in the presence of high-dimensional data with intrinsic low-rank structure. In this paper, we present a simple yet powerful method to solve an as yet unaddressed core problem for the hypergeometric distribution, namely inferring a mixture of discrete distributions where both the size of the total population and its constituent categories are unknown. We first show that parameter inference is tractable when there are two or more categories using empirical simulations, and that our method outperforms existing likelihood-based approaches. To demonstrate its value in real-world settings, we use our method to infer a latent vocabulary for reading passages, which can be used to assess their complexity. Finally, we address the technical limitations of genomics experimental methods and recover missing gene transcript counts in high-throughput single-cell genomics measurements.
@@ 2.1 The Hypergeometric Distribution
Consider an urn that contains $N$ balls divided into $K=2$ categories (e.g. colors): $N_{1}$ balls are white and $N_{2}=N-N_{1}$ are black. Each ball has equal probability of being selected from the urn. If we sample $n<N$ balls without replacement, obtaining $c_{1}$ white ball counts and $c_{2}$ black counts, the distribution of the number of sampled balls of each color is given by the hypergeometric distribution, whose joint probability mass function is (Moivre, 1711):
@@
Typically $N$ is assumed to be known, however here we consider the setting where neither $N$ nor any $N_{i}$ are known. In the typical setting the distribution for $K=2$ is called the univariate hypergeometric distribution, with $K>2$ referred to as multivariate, but because in our problem setting there are already two unknown variables when $K=2$, we use the term hypergeometric distribution for any $K\geq 2$.
@@ 2.2 Existing Maximum Likelihood Estimators
Known total population size: When the total population size $N$ is known, the object is to estimate the true number of elements $N_{i}$ in each constituent category $i\in 1,\ldots,K$. The maximum likelihood estimator is then essentially the known total population size scaled by the sample frequency of each category, with adjustment to ensure a correct integer solution (Oberhofer & Kaufmann, 1987).
Unknown total population size: In a more complex case known as the capture-recapture problem (Darroch, 1958), the total population size $N$ is unknown. To estimate the total population, a sample is first drawn from the underlying distribution. All objects belonging to one of the categories in this sample are tagged, and all sampled elements are returned to the population. A second sample from the same distribution is then drawn, and the number of tagged samples that reappear permits the estimation of the total population size using maximum likelihood. This method has found important applications in biology and ecology. However, it depends on the ability to tag and resample the same population, which is often not possible in practice, and does not extend to a mixture of distributions.
Additionally, Tohma et al. (1991) propose a number of approximate methods for the estimation of the parameters of the hypergeometric distribution when the category sizes are unknown, including an approximation using the likelihood, however they do not maximize the likelihood directly.
@@ 2.3 Related Methods
The multinomial distribution, which is the limiting form of the hypergeometric distribution when the sample size is negligible relative to the population ($N>>n$), appears frequently in machine learning literature. For example, it is often used to model counts based on their relative abundance. Latent Dirichlet allocation (Blei et al., 2003) uses a hierarchical generative framework to model the distribution of topics, documents over topics, and word (category) counts from a vocabulary (population) over documents. The distributions over topics and words are multinomial, with word counts being transformed into their relative frequency in the document. (Liang et al., 2018) use a variational autoencoder with multinomial likelihood applied to collaborative filtering for recommender systems. The click data and play counts are binarized to accommodate the multinomial likelihood.
@@ 3.1 Single Ground-Truth Distribution
We first assume that there exists a single true discrete population of elements, with total number of elements $N$ belonging to $K$ categories. We consider the setting where none of the category sizes are known, and we wish to estimate them from the data; that is, we do not know any category size $N_{i}$, nor the total population size $N=\sum_{i=1}^{K}N_{i}$. We assume under-sampling of the underlying distribution ($n<<N$) as this is typical real-world applications.
@@
$\mathcal{L}(N_{1},\ldots,N_{K}|{\bm{c}}_{1},\ldots,{\bm{c}}_{T})=\prod_{t=1}^{T}\frac{\prod_{i=1}^{K}\binom{N_{i}}{c_{t,i}}}{\binom{\sum_{i=1}^{K}N_{i}}{\sum_{i=1}^{K}c_{t,i}}}$ (3)
@@
$\begin{split}&\log\mathcal{L}(N_{1},\ldots,N_{K}|{\bm{c}}_{1},\ldots,{\bm{c}}_{T})=\\ &\sum_{t=1}^{T}\left[\sum_{i=1}^{K}\log\binom{N_{i}}{c_{t,i}}-\log\binom{\sum_{i=1}^{K}N_{i}}{\sum_{i=1}^{K}c_{t,i}}\right]\end{split}$ (4)
The hypergeometric distribution is not part of the exponential family, so it is not obvious that a closed-form maximum likelihood estimator exists, and therefore we turn to numerical optimization methods. To enable continuous optimization, we consider a continuous and differentiable relaxation of the log-likelihood by replacing the factorials in the binomial coefficient with the gamma function, which is the extension of the factorial to real-numbered arguments:
@@
Using the resulting log-likelihood for the parameter set $\theta=\{N_{1},\ldots,N_{K}\}$, we perform maximum likelihood estimation to obtain the MLE $\hat{\theta}\equiv\{\hat{N}_{1},\ldots,\hat{N}_{K}\}=\operatorname*{arg\,max}_{\theta}\log\mathcal{L}(\theta|{\bm{c}}_{1},\ldots,{\bm{c}}_{T})$. During training, the continuous estimates of $N_{i}$ are used to perform gradient updates, whereas during inference the estimates can be rounded to the nearest integer if a discrete solution is required.
The binomial coefficient is not defined if $c_{k}>\hat{N}_{k}$, which would correspond to the impossible scenario of sampling more balls of a given color than are present in the urn. To impose the requirement $\hat{N}_{k}\geq c_{k}$, we add a violation penalty $C_{\text{viol}}$ to the negative log-likelihood we seek to minimize (Equation 6), and we threshold any estimates of $\hat{N}_{k}<c_{k}$ at $c_{k}$ before evaluating the likelihood (Equation 7). The use of a violation function is a common approach to solving constrained optimization problems (Bertsekas, 2014). We threshold $N_{i}$ at the observed sample value $c_{t,i}$ as opposed to the minimum across all samples $\min_{t}c_{t,i}$ to remain as general as possible. This is because in the case of a mixture of distributions we do not know which observation originates from which underlying distribution, and hence do not know what the correct minimum is.
$C_{\text{viol}}$ $=\sum_{i=1}^{K}\max(0,c_{i}-\hat{N}_{i})$ (6)
@@
In this paper we are interested in modelling scenarios where we have access to many samples drawn from the same underlying population, but where this population is under-sampled. Specifically, if $N$ is the total population size, we assume that we obtain samples with at most $n_{max}$ objects drawn from the ground-truth distribution in each trial, giving a sample fraction $f_{max}=n_{max}/N$. For example, if $f_{max}=0.5$ and the true distribution has 100 elements, no observation will contain more than 50 samples.
@@ 3.2 Mixture of Ground-Truth Distributions
While the hypergeometric likelihood can be used to directly estimate the ground truth for a single underlying distribution using maximum likelihood (Section 4.1), data generated by real world processes is often better represented by a mixture of distributions. Therefore, we next propose a generative latent variable model $p({\bm{c}},{\bm{z}})=p({\bm{c}}|{\bm{z}})p({\bm{z}})$ to jointly model the observed counts conditional on a continuous latent variable ${\bm{z}}$.
We extend our approach to allow for the estimation of a high-dimensional distribution that is conditional on a continuous latent variable, allowing it to capture a continuous mixture of count distributions. This modelling assumption is essential when the true distribution is considerably undersampled, leading to sparsity, and so information needs to be shared between similar observations to successfully model the data. The variational autoencoder (VAE) (Kingma & Welling, 2014) is a powerful framework for performing efficient estimation of the generative distribution parameters in the presence of a continuous latent variable and high-dimensional data.
@@
$\begin{split}\mathcal{L}(\psi,\phi;{\bm{c}}_{t})=&-D_{\text{KL}}(q_{\phi}({\bm{z}}|{\bm{c}}_{t})||p({\bm{z}}))+\mathbb{E}_{q_{\phi}}\left[\log p_{\psi}({\bm{c}}_{t}|{\bm{z}})\right]\\ &+\sum_{i=1}^{K}\max(0,c_{i}-\hat{N}_{i})\end{split}$ (8)
The first term is the KL divergence between the approximate posterior and the prior, the second term is the expected log-likelihood under the approximate posterior, and the third term is the violation penalty. Optimization of this objective enables the estimation of counts under-sampled from a mixture of distributions. The MLE estimate for $\theta_{i}$ is obtained by first computing ${\bm{z}}_{i}=f({\bm{c}}_{i})$, where ${\bm{c}}_{i}$ is the observed count vector, then computing $\theta_{i}=g({\bm{z}}_{i})$, where the output layer of $g$ is a linear layer with $K$ features and a ReLU activation.
@@ 4.1 Tractability of Inferring Ground-Truth Distributions
Using empirical data simulation, we begin by demonstrating that this inference problem is tractable in the ideal case where observations are sampled from a single discrete ground-truth distribution. Despite its simplicity, this scenario has yet to have been shown tractable, and in fact appears to have generally been assumed to be intractable.
@@
To demonstrate the typical behavior, we show results for a simulation for the scenario where $N_{1}=70$ and $N_{2}=30$ ($N=100$, $K=2$, $f_{max}=0.4$), evaluating the negative log-likelihood (NLL) for all possible combinations of $\hat{N}_{1}$ and $\hat{N}_{2}$. Figure 1 shows the resulting loss landscape, and we see that the minimum NLL value corresponds to the true $N_{1}$ and $N_{2}$ values. The lowest NLL region lies along the line of correct $N_{2}/N_{1}$ ratio. We note that the difference in NLL between the optimum and surrounding estimates is small, especially for estimates with the same ratio $N_{2}/N_{1}$.
@@
We next investigate the robustness of the maximum likelihood estimate for a wide range of numbers of observations (trials) $T$ and max sample fractions $f_{max}$. As the data is discrete, we use the Manhattan distance between the true and estimated counts to quantify the estimation error. In Figure 2, we see that the location of the minimum NLL converges to the true value as the number of observations increases, for different levels of under-sampling. As expected, a higher max sample fraction results in faster convergence and more accurate estimates, as quantified by lower error for the same number of samples. We extend this experiment to $K=3$ (Figure 7, Appendix E), and find that increasing the number of categories reduces the number of observations required to reach zero error. Here we use a confidence interval of 50% for visual clarity, and the same figure with a 90% confidence interval is included in Appendix E.
Figure 2: Maximum likelihood estimate Manhattan error for different numbers of trials at different max sample fractions. 50% confidence interval over 50 random seeds. $K=2,N=100$ $(N_{1}=40,N_{2}=60)$